I would greatly appreciate your help! White pebbles are sold for $4.00 lb. and reddish pebbles are sold for $2.00 lb.. If 75 lbs. of white pebbles were combined with 25 lbs. of reddish pebbles, who much would the price per lb. be for the 100 lbs. of pebbles? Thanks!
Answer by checkley77(12844)

75*4+25*2=100X
300+50=100X
100X=350
X=350/100
X=$3.50 PRICES OF THE MIXED PEBBLES PER LB.